0 0
Add a comment Improve this question Transcribed image text
Answer #1
Mass is constant and is kept as M.
car1: F=ma
10=ma
a=10/m

Car2:
Net force = 10-(2*5/√2)
               =2.93
a=2.93/m

Car3:all forces are in vertical, so no horiz acc.
a=0

Car4:F=0, ∴a=0;

Car5:Net force =10/√2-5
                       =2.07 backwards
a=-2.07/m (deceleration)

Car6:net force =10-10 =0
a=0;

Therefore after a small time period say Δt, the car with themost acceleration will have the most speed.
ranking them(largest to smallest):
1,2,(3,4,6),5 (where 3,4, and 6 have the samevelocity)